OpenStudy (anonymous):

Hydrochloric acid reacts with calcium to form hydrogen and calcium chloride. If 100 grams of hydrochloric acid reacts with 100 grams of calcium, what is the limiting reactant? 2HCl + Ca → CaCl2 + H2 calcium hydrogen calcium chloride hydrochloric acid

OpenStudy (anonymous):

medals to whoever answers first need answer fast

OpenStudy (rushwr):

What do u think the answer is ?

OpenStudy (rushwr):

First of all we can cut off cacl2 and H2 because they aren't reactants. Those are the products. So we are left with Ca and HCl then we have to find the no. of moles of each reactants moles = mass divided by molar mass molar mass of HCl = 36.5gmol^-1 Molar mass of Ca = 40gmol^-1 We know the masses 100g each find the moles of each Moles of Ca = 100/40 = 2.5moles moles of Ca = 100/36.5 = 2.7 moles Anyways according to the reaction u need more HCl to make sure than u finish Ca. The stoicheometry of HCl : Ca = 2:1 So the limitting agent is the insufficient one that is HCl